OpenStudy (anonymous):

A group of friends were voting for their favorite song: two-fifths chose song A, one-sixth of the remaining friends chose song B, and the rest chose song C. If there were thirty friends in all, how many chose song C? ten twelve fifteen eighteen

OpenStudy (triciaal):

A + B + C = 30 A = 2/5*30 = 12 B= 3/5*1/6*30 = 3 C= 30*3/5*5/6 = 15